OpenStudy (anonymous):

5/3x+4=2/3x solve for X.

OpenStudy (anonymous):

5/3x+4=2/3x 5/3x - 2/3 x = -4 3/3x = -4 x = -4

OpenStudy (anonymous):

x=-1/4

OpenStudy (anonymous):

I got positive 4, thank you, I see my mistake now

OpenStudy (anonymous):

you are welcome :)

OpenStudy (anonymous):

is -1/4 acceptable?

OpenStudy (anonymous):

that's the answer i got i used an app and 100% of the time it's correct it's never gave me an wrong answer